OpenStudy (anonymous):

The point (-3/5,y) in the third quadrant corresponds to angle θ on the unit circle. The value of sec θ is? and the value of cot θ is?

OpenStudy (anonymous):

@jameskang1015

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

-3/5 = cos Θ because the x-coordinate is cosine. Flip it over to get sec Θ

OpenStudy (anonymous):

so would sec 0 be -5/3? confused

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

\(sin^2 \theta + cos^2 \theta=1\) where sin Θ = y \[y^2+ \left( \frac{ -3 }{ 5 } \right)^2=1\] take y to be the negative solution since Θ is in the third quadrant y = -4/5 cot Θ = x/y

OpenStudy (anonymous):

\[\cot\theta=-\frac{ 3 }{ 5 }\div -\frac{ 4 }{ 5 }\]